# On Modular Invariants of Truncated Polynomial Rings in low ranks

### v1

@@ 2.5. The Borel subgroup
In [6], we have constructed a basis for the invariant ring of ${\mathcal{Q}}_{m}^{B}$ under the action of the Borel subgroup for arbirary $m$ and $n$. In the case $n=3$, the $\mathbb{F}_{q}$-basis $\mathcal{B}_{m}(3)$ for ${\mathcal{Q}}_{m}^{B}$ consists 4 families:
@@ 3. A lower bound on the total dimension of the invariant subspaces
The original parabolic conjecture predicted that the Hilbert-Poincaré series for the space of invariants ${\mathcal{Q}}_{m}^{\alpha}(n)$, where $\alpha$ is a composition of $n$, is the (finite) polynomial $C_{\alpha,m}(t)$, whose explicit form is given in 1.1. In particular, the value of $C_{\alpha,m}(t)$ when $t=1$ should be the total dimension of the graded $\mathbb{F}_{q}$-vector space ${\mathcal{Q}}_{m}^{\alpha}(n)$.
@@
For each $m,n\geq 1$ and any composition $\alpha$ of $n$, the total dimension of the graded vector space ${\mathcal{Q}}_{m}^{\alpha}(n)$ is at least $C_{\alpha,m}(1)$.
This result is already implicit in the discussion in Section 6 of [10], and is valid for any $n$. This relatively simple remark has the following important consequence in practice: suppose we can construct a generating set for the invariant subspace ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$ such that, by dimension counting, the Hilbert-Poincare series $C^{\prime}_{m,\alpha}(t)$ for ${\mathcal{Q}}_{m}^{\alpha}(n)$ is not greater than $C_{\alpha,m}(t)$ ($f(t)\leq g(t)$ iff $g-f$ is a polynomial with non-negative coefficients). Then since there is a reverse inequality $C^{\prime}_{\alpha,m}(1)\geq C_{\alpha,m}(1)$ when evaluated at $t=1$, we can conclude that these two series are identical since both are finite degree polynomials with non-negative coefficients. Hence our generating set is actually a basis for ${\mathcal{Q}}_{m}^{P(\alpha)}(n)$.
Proof of Proposition 3.1: Consider the following ungraded quotient of the polynomial algebra
@@
The action of $\operatorname{GL}_{n}$ on $\operatorname{\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]$ descends to this quotient. There is a natural filtration:
@@
then induces an isomorphism of $\operatorname{GL}_{n}$-equivariant $\operatorname{\mathbb{F}}_{q}$-algebras:
@@
On the other hand, if we apply the functor taking $\operatorname{GL}_{n}$-invariant on each short exact sequence of $\operatorname{GL}_{n}$-modules
@@
and computing the dimension, we obtain an inequality
$\dim(F_{i})^{\operatorname{GL}_{n}}\leq\dim(F_{i-1})^{\operatorname{GL}_{n}}+\dim(F_{i}/F_{i-1})^{\operatorname{GL}_{n}}$
@@
$\dim{\mathcal{R}}^{\operatorname{GL}_{n}}\leq\dim(\operatorname{\mathfrak{gr}}_{F}{\mathcal{R}})^{\operatorname{GL}_{n}}\cong\dim{\mathcal{Q}}_{m}^{\operatorname{GL}_{n}}(n).$
To compute the $\operatorname{\mathbb{F}}_{q}$-dimension of ${\mathcal{R}}^{\operatorname{GL}_{n}}$, we extend the field $\operatorname{\mathbb{F}}_{q}$ to a field $\operatorname{\mathbb{F}}$ which contains $\operatorname{\mathbb{F}}_{q^{m}}$. Then the evaluation map
@@
This isomorphism is also $\operatorname{GL}_{n}$-equivariant where $\operatorname{GL}_{n}$ acts on the right-hand side via the embedding $\operatorname{GL}_{n}=\operatorname{GL}_{n}(\operatorname{\mathbb{F}}_{q})\subset\operatorname{GL}_{n}(\operatorname{\mathbb{F}}_{q^{m}}).$ It follows that there is an isomorphism
$(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}})^{\operatorname{GL}_{n}}\cong\operatorname{Map}\big{(}\operatorname{\mathbb{F}}_{q^{m}}^{n}/\operatorname{GL}_{n},\operatorname{\mathbb{F}}\big{)}.$
The $\operatorname{\mathbb{F}}$-dimension of $(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}})^{\operatorname{GL}_{n}}$ is thus equal to the cardinality of the orbit set $(\operatorname{\mathbb{F}}_{q^{m}})^{n}/\operatorname{GL}_{n}$. Finally this cardinality is equal to the sum $\sum_{s=0}^{\min(m,n)}\begin{bmatrix}m\\ s\end{bmatrix}_{q}$ where $\begin{bmatrix}m\\ s\end{bmatrix}_{q}$ is the $q$-binomial coefficient which counts the number of $s$-dimensional subspaces of the $\operatorname{\mathbb{F}}_{q}$-vector space $\operatorname{\mathbb{F}}_{q^{m}}$.
@@
$\dim_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}}^{\operatorname{GL}_{n}}=\dim_{F}(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}}^{\operatorname{GL}_{n}})=\sum_{s=0}^{\min(m,n)}\begin{bmatrix}m\\ s\end{bmatrix}_{q}=C_{m,n}(1)\leq\dim{\mathcal{Q}}_{m}^{\operatorname{GL}_{n}}(n).$
The argument above clearly is not affected if one uses a parabolic subgroup $P_{\alpha}$ instead of $\operatorname{GL}_{n}$. The proposition is proved. ∎
@@
For each composition $\alpha$ of $n$, the Hilbert-Poincaré series for the $\mathbb{F}_{q}$-space spanned by $\mathcal{B}_{m}(\alpha)$ is not greater than $C_{\alpha,m}(t)$.
@@
Recall that
@@
where
$e(m,\alpha,\beta)=\sum(\alpha_{i}-\beta_{i})(q^{m}-q^{B_{i}})=(\alpha_{1}-\beta_{1})(q^{m}-q^{\beta_{1}})+q^{\beta_{1}}\sum_{i\geq 2}(\alpha_{i}-\beta_{i})(q^{m-\beta_{1}}-q^{\beta_{2}+\ldots+\beta_{i}}).$
Also, from [15, Section 7], we have the following identity for the $(q,t)$-multinomial coefficient:
$\begin{bmatrix}m\\ \beta,m-|\beta|\end{bmatrix}_{q,t}=\begin{bmatrix}m\\ \beta_{1}\end{bmatrix}_{q,t}\;\Biggl{(}\begin{bmatrix}m-\beta_{1}\\ \beta_{2},\ldots,\beta_{\ell},m-|\beta|\end{bmatrix}_{q,t}\biggl{)}^{q^{\beta_{1}}}.$
Thus if we write $\beta^{\prime}=(\beta_{2},\ldots,\beta_{\ell})$ and similarly for $\alpha^{\prime}$, then $C_{\alpha,m}(t)$ can be written as
$\sum_{\beta_{1}\leq\min(\alpha_{1},m)}t^{(\alpha_{1}-\beta_{1})(q^{m}-q^{\beta_{1}})}\begin{bmatrix}m\\ \beta_{1}\end{bmatrix}_{q,t}\biggl{(}\sum_{\beta^{\prime}\leq\alpha^{\prime},|\beta^{\prime}|\leq m-\beta_{1}}t^{e(m-\beta_{1},\beta^{\prime},\alpha^{\prime})}\begin{bmatrix}m-\beta_{1}\\ \beta^{\prime},m-\beta_{1}-|\beta^{\prime}|\end{bmatrix}_{q,t}\biggl{)}^{q^{\beta_{1}}}.$
Comparing this last formula with the recursive definition of $\mathcal{B}_{m}(\alpha)$, noting that $\delta_{s+1;m}^{\alpha_{1}-s}$ raises degree by $(\alpha_{1}-s)(q^{m}-q^{s})$, the Frobenius operator multiply the degree by $q$, and the Hilbert-Poincaré series for $\Delta^{m}_{s}$ is exactly $\begin{bmatrix}m\\ s\end{bmatrix}_{q,t}$, we see that the result follow easily by induction. ∎
The results in this section simplify our tasks considerably. Indeed, in order to show that $\mathcal{B}_{m}(\alpha)$ is a basis for ${\mathcal{Q}}_{m}^{P_{\alpha}}(n)$, it suffices to verify that it is a generating set.
@@
As remarked in the previous section, it suffices to show that $\mathcal{B}_{m}(2)$ is a generating set for ${\mathcal{Q}}_{m}^{G}(2)$. The proof of the Proposition 1.6 is done in two steps. Firstly, we use the relative transfer from the known invariant ring ${\mathcal{Q}}^{B}$ of the Borel subgroup $B$ to show that a bigger set $\mathcal{B}^{\prime}$, consisting of 3 families:
@@
is a generating set for ${\mathcal{Q}}_{m}^{G}(2)$. Then we show that to generate the invariant ring, it is possible to replace the Dickson algebra ${\mathcal{D}}_{2}$ in the third family by the smaller subspace $\Delta^{m}_{2}$.
@@
The set $\mathcal{B}^{\prime}$ spans ${\mathcal{Q}}_{m}^{G}(2)$.
@@
In [6], we showed that an $\mathbb{F}_{q}$-basis $\mathcal{B}_{m}(1,1)$ for the ring of invariants under the Borel subgroup consisting of the following 2 families
@@
It is a standard fact about the relative transfer $\operatorname{Tr}\colon{\mathcal{Q}}_{m}^{B}(2)\to{\mathcal{Q}}_{m}^{G}(2)$ that it is onto, so that a generating set for ${\mathcal{Q}}^{G}(2)$ can be obtained from the image of $\operatorname{Tr}$.
By direct inspection, we have the following identity in ${\mathcal{Q}}_{m}$ for any $0\leq s<[m]_{q}$:
@@
Since $y_{s}$ is already a $\operatorname{GL}_{2}$-invariant, this formula allows us to compute the transfer of the first family:
@@
Clearly, $\operatorname{Tr}(x_{1}^{(s+1)(q-1)})$ is again a Dickson polynomial in ${\mathcal{D}}_{2}$. By Proposition 2.5, the product of $y_{0}$ with a Dickson polynomial is either zero or another element in the $y$ family. Thus $\operatorname{Tr}(x_{1}^{q^{m}-1}x_{2}^{a(q-1)})$ is either $-y_{s+1}$ or zero.
The transfer of the second family $Q_{1,0}^{i_{1}}Q_{2,1}^{i_{2}}$ is clearly a Dickson polynomial. The lemma follows. ∎
@@
For each $10\leq i\leq m-$ 1, we have the following decomposition in ${\mathcal{Q}}_{m}(2)$:
@@
It is easy to see, by determinant manipulation, that $\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})=y_{\frac{q^{i}-1}{q-1}}$ is a genuine Dickson polynomial. When $i=0$, the Dickson monomial $y_{0}$ must contains $Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}$ as a montrivial summand since both are reduced to $x_{1}^{q^{m}-q}$ when setting $x_{2}=0$. Furthermore, $y_{0}-Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}$ is divisible by $Q_{2,0}$. Note that in degree $q^{m}-q$, $Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}$ is the only non-essential Dickson monomial. We have proved that
@@
It is easy to check that if $i>[m-1]_{q}$, then $Q_{2,0}^{i}=L_{2}^{i(q-1)}=0$ in ${\mathcal{Q}}_{m}(2)$. Also, by induction on $m$, we have another amusing formula:
@@
It is straightforward to check that all elements in the six families listed in Proposition 1.7 are $P(2,1)$-invariants. It suffices to show that $\mathcal{B}_{m}(2,1)$ is a generating set for ${\mathcal{Q}}_{m}^{P(2,1)}(3)$.
Suppose $f$ is a nonzero polynomials in ${\mathcal{Q}}_{m}^{P(2,1)}(3)$. If $f$ is considered as a polynomial in $x_{3}$, then its coefficients are $\operatorname{GL}_{2}$-invariant in the $2$ variables $x_{1},x_{2}$, which are classified in Section 4. If one of these coefficients has the form $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$, then the corresponding term belongs to the last 2 families in the list. By subtracting it if necessary, let us assume that this term does not occur in $f$.
Now suppose that the $x_{1}$ (and hence $x_{2}$ by symmetry) degree of $f$ is strictly less than $q^{m}-1$. Then $f$ is a linear combination of polynomials of the last two types in the list of Borel invariants in 2.5. In particular, the $x_{3}$-degree of $f$ must be divisible by $q^{2}-q$. Write $f$ as
$f=x_{3}^{aq(q-1)}f_{aq}+x_{3}^{(aq-1)(q-1)}f_{aq-1}+\ldots,$
where $f_{i}=f_{i}(x_{1},x_{2})\in{\mathcal{Q}}_{m}^{G}(2)$. If $f_{aq}$ contains $y_{t}$ for some $t\geq 0$, then we can subtract from $f$ an appropriate polynomial in the third or fourth family in the list 1.7. So we can assume that $f_{aq}$ belongs to $\Delta^{m}_{2}$. But this implies that this top degree monomial is a non-trivial summand of a $B$-invariant polynomial of the form $x_{1}^{j_{1}(q-1)}Q_{2,1}^{j_{2}}Q_{3,2}^{j_{3}}$. In particular, this shows that $a$ is divisible by $q$. Write $a=bq$. Then by subtracting from $f$ an appropriate linear combination of first two families of $\mathcal{B}_{m}(2,1)$, we will obtain a $G(2,1)$-invariant polynomial of strictly smaller $x_{3}$-degree. The result follows by induction on $x_{3}$-degree. ∎
@@
We can also prove that $\mathcal{B}_{m}(2,1)$ is linearly independent as follows. Consider these elements of $\mathcal{B}_{m}(2,1)$ as polynomials in $x_{3}$. The table below records the coefficient of the highest $x_{3}$-exponent.
$\mathcal{B}_{m}(2,1)$ $x_{3}$-degree Coefficient
@@
It suffices to show that $\mathcal{B}_{m}(1,2)$ is a generating set, and the proof is by downward induction on the smallest $x_{1}$-degree. Write

### v2

@@ 2.5. The Borel subgroup
In [6], we constructed a basis for the invariant ring of ${\mathcal{Q}}_{m}(n)$ under the action of the Borel subgroup for arbitrary $m$ and $n$. In the case $n=3$, the $\mathbb{F}_{q}$-basis $\mathcal{B}_{m}(3)$ for ${\mathcal{Q}}_{m}^{B}$ consists of $4$ families:
@@ 3. A lower bound on the total dimension of the invariant subspaces
The original Parabolic Conjecture predicted that the Hilbert series of the invariant space ${\mathcal{Q}}_{m}(n)^{P(\alpha)}$, where $\alpha$ is a composition of $n$, is given by the (finite) polynomial $C_{\alpha,m}(t)$, whose explicit form is stated in Conjecture 1.1. In particular, the value of $C_{\alpha,m}(t)$ when $t=1$ is expected to equal the total dimension of the finite-dimensional graded $\mathbb{F}_{q}$-vector space ${\mathcal{Q}}_{m}(n)^{P(\alpha)}$.
@@
For each $m,n\geq 1$ and any composition $\alpha$ of $n$, the total dimension of the graded vector space ${\mathcal{Q}}_{m}(n)^{P(\alpha)}$ is at least $C_{\alpha,m}(1)$.
This result is already implicit in [9, Section 6] and holds for arbitrary $n\geq 1$. This relatively simple observation has an important practical consequence: suppose we can construct a generating set for the invariant subspace ${\mathcal{Q}}_{m}(n)^{P(\alpha)}$ such that, by dimension counting, the associated Hilbert series $C^{\prime}_{m,\alpha}(t)$ does not exceed $C_{\alpha,m}(t)$ (i.e. $f(t)\leq g(t)$ if and only if $g(t)-f(t)$ is a polynomial with non-negative coefficients). Then, since the reverse inequality $C^{\prime}_{\alpha,m}(1)\geq C_{\alpha,m}(1)$ holds evaluated at $t=1$, we may conclude that these two series must be identical since both are finite degree polynomials with non-negative coefficients. It follows that our generating set is, in fact, a basis for ${\mathcal{Q}}_{m}(n)^{P(\alpha)}$.
Proof of Proposition 3.1: Consider the following ungraded quotient of the polynomial algebra:
@@
The action of $G=GL_{n}$ on $\operatorname{\mathbb{F}}_{q}[x_{1},\ldots,x_{n}]$ descends to this quotient. There is a natural filtration:
@@
induces an isomorphism of $G$-equivariant $\operatorname{\mathbb{F}}_{q}$-algebras:
@@
On the other hand, if we apply the functor taking $G$-invariants on each short exact sequence of $G$-modules
@@
and compute the dimension, we obtain an inequality
$\dim(F_{i})^{G}\leq\dim(F_{i-1})^{G}+\dim(F_{i}/F_{i-1})^{G}$
@@
$\dim{\mathcal{R}}^{G}\leq\dim(\operatorname{\mathfrak{gr}}_{F}{\mathcal{R}})^{G}=\dim{\mathcal{Q}}_{m}(n)^{G}.$
To compute the $\operatorname{\mathbb{F}}_{q}$-dimension of ${\mathcal{R}}^{G}$, we extend the field $\operatorname{\mathbb{F}}_{q}$ to a field $\operatorname{\mathbb{F}}$ which contains $\operatorname{\mathbb{F}}_{q^{m}}$. Then the evaluation map
@@
This isomorphism is also $G$-equivariant where $G$ acts on the right-hand side via the embedding $G=GL_{n}(\operatorname{\mathbb{F}}_{q})\subset GL_{n}(\operatorname{\mathbb{F}}_{q^{m}}).$ It follows that there is an isomorphism
$(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}})^{G}\cong\operatorname{Map}\big{(}\operatorname{\mathbb{F}}_{q^{m}}^{n}/G,\operatorname{\mathbb{F}}\big{)}.$
The $\operatorname{\mathbb{F}}_{q}$-dimension of $(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}})^{G}$ is thus equal to the cardinality of the orbit set $\operatorname{\mathbb{F}}_{q^{m}}^{n}/G$. Finally this cardinality is equal to the sum $\sum_{s=0}^{\min(m,n)}\begin{bmatrix}m\\ s\end{bmatrix}_{q}$ where $\begin{bmatrix}m\\ s\end{bmatrix}_{q}$ is the $q$-binomial coefficient which counts the number of $s$-dimensional subspaces of the $\operatorname{\mathbb{F}}_{q}$-vector space $\operatorname{\mathbb{F}}_{q^{m}}$.
@@
$\dim_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}}^{G}=\dim_{F}(\operatorname{\mathbb{F}}\otimes_{\operatorname{\mathbb{F}}_{q}}{\mathcal{R}}^{G})=\sum_{s=0}^{\min(m,n)}\begin{bmatrix}m\\ s\end{bmatrix}_{q}=C_{m,n}(1)\leq\dim{\mathcal{Q}}_{m}(n)^{G}.$
The argument above clearly is not affected if one uses a parabolic subgroup $P_{\alpha}$ instead of $G$. The proposition is proved. ∎
@@
For each composition $\alpha$ of $n$, the Hilbert series for the $\mathbb{F}_{q}$-space spanned by $\mathcal{B}_{m}(\alpha)$ is not greater than $C_{\alpha,m}(t)$.
@@
This is almost obvious from the definition of $C_{\alpha,m}(t)$ and the relationship between the subset $\Delta^{m}_{s}\subset{\mathcal{D}}_{s}$ and the (multinomial) $(q,t)$-coefficient. Recall that
@@
where $e(m,\alpha,\beta)=\sum(\alpha_{i}-\beta_{i})(q^{m}-q^{B_{i}})$. Our conjectural basis explains precisely how the summands in the above expression occur. Indeed, for each composition $\beta$, with $\beta\leq\alpha$ and $|\beta|\leq m$, $e(m,\alpha,\beta)$ is exactly the total degree raised by the delta operators in a basis element
$\delta_{B_{1}+1;m}^{\alpha_{1}-\beta_{1}}\bigg{(}f_{1}\delta_{B_{2}+1;m}^{\alpha_{2}-\beta_{2}}\big{(}\cdots f_{\ell-1}\delta_{B_{\ell}+1;m}^{\alpha_{\ell}-\beta_{\ell}}(f_{\ell})\big{)}\bigg{)},$
and the $(q,t)$-multinomial coefficient $\begin{bmatrix}m\\ \beta,m-|\beta|\end{bmatrix}_{q,t}$ corresponds to the number of choices of $\ell$-tuple $(f_{1},\ldots,f_{\ell})$ where $f_{i}\in\Phi^{B_{i-1}}\Delta_{\beta_{i}}^{m-B_{i-1}}$. Note that according to [12, (7.1)], we can write
$\begin{bmatrix}m\\ \beta,m-|\beta|\end{bmatrix}_{q,t}=\begin{bmatrix}m\\ \beta_{1}\end{bmatrix}_{q,t}\varphi^{\beta_{1}}\begin{bmatrix}m-B_{1}\\ \beta_{2}\end{bmatrix}_{q,t}\varphi^{\beta_{2}}\begin{bmatrix}m-B_{2}\\ \beta_{3}\end{bmatrix}_{q,t}\ldots$
where $\varphi$ is the (genuine) Frobenius operator, which reflects how our operator $\Phi$ affects the degree. ∎
The results in this section simplify our tasks considerably. Indeed, in order to show that $\mathcal{B}_{m}(\alpha)$ is a basis for ${\mathcal{Q}}_{m}(n)^{P_{\alpha}}$, it suffices to verify that it is a generating set.
@@
As remarked in the previous section, it suffices to show that $\mathcal{B}_{m}(2)$ is a generating set for ${\mathcal{Q}}_{m}(2)^{G}$. The proof of Proposition 1.6 is done in two steps. Firstly, we use the relative transfer from the known invariant ring ${\mathcal{Q}}^{B}$ of the Borel subgroup $B$ to show that a bigger set $\mathcal{B}^{\prime}$, consisting of 3 families:
@@
is a generating set for ${\mathcal{Q}}_{m}(2)^{G}$. Then we show that to generate the invariant ring, it is possible to replace the Dickson algebra ${\mathcal{D}}_{2}$ in the third family by the smaller subspace $\Delta^{m}_{2}$.
@@
The set $\mathcal{B}^{\prime}$ consisting of 3 families above spans ${\mathcal{Q}}_{m}(2)^{G}$.
@@
In [6], we described an $\mathbb{F}_{q}$-basis $\mathcal{B}_{m}(1,1)$ for the ring of invariants under the Borel subgroup consisting of the following 2 families:
@@
It is a standard fact about the relative transfer $\operatorname{Tr}\colon{\mathcal{Q}}_{m}(2)^{B}\to{\mathcal{Q}}_{m}(2)^{G}$ that it is onto since the index of $B$ in $G$ is relatively prime to $p$ (see [14, Section 2.2]), so a generating set for ${\mathcal{Q}}_{m}(2)^{G}$ can be obtained from the image of $\operatorname{Tr}$.
In what follows, we adopt Goyal’s notation $y_{s}$ for $\delta_{2}(Q_{1,0}^{s})$ for the sake of brevity. By direct inspection, we have the following identity in ${\mathcal{Q}}_{m}(2)$ for any $0\leq s<[m]_{q}$:
@@
Since $y_{s}$ is already a $G_{2}$-invariant, this formula allows us to compute the transfer of the first family:
@@
Note that $\operatorname{Tr}(x_{1}^{(s+1)(q-1)})$ is a Dickson polynomial divisible by $Q_{2,0}$ because $x_{1}^{(s+1)(q-1)}$ is a genuine $B$-invariant which vanishes when setting $x_{1}=0$. By the first two parts of Proposition 2.7, the product of $y_{0}$ with $Q_{2,0}$ is trivial in ${\mathcal{Q}}_{m}(2)$. Thus
$\operatorname{Tr}(x_{1}^{q^{m}-1}x_{2}^{s(q-1)})=-y_{s+1}.$
For the second family, we argue similarly, noting that
$\operatorname{Tr}(Q_{1,0}^{i_{1}}Q_{2,1}^{i_{2}})=\operatorname{Tr}(Q_{1,0}^{i_{1}})Q_{2,1}^{i_{2}}\in{\mathcal{D}}_{2}.$
The lemma follows. ∎
@@
For each $10\leq i\leq m-$ 1, we have the following decomposition in $S$:
@@
It is easy to see, by determinant manipulation, that $\delta_{2}(Q_{1,0}^{\frac{q^{i}-1}{q-1}})=y_{\frac{q^{i}-1}{q-1}}$ is a genuine Dickson polynomial. When $i=0$, the Dickson monomial $y_{0}$ must contain $Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}$ as a nontrivial summand since both are reduced to $x_{1}^{q^{m}-q}$ when setting $x_{2}=0$. Furthermore, $y_{0}-Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}$ is divisible by $Q_{2,0}$. Note that in degree $q^{m}-q$, $Q_{2,1}^{\frac{q^{m-1}-1}{q-1}}$ is the only non-essential Dickson monomial. We have proved that
@@
It is easy to check that if $i>[m-1]_{q}$, then $Q_{2,0}^{i}=L_{2}^{i(q-1)\geq q^{m-1}}=0$ in ${\mathcal{Q}}_{m}(2)$. Also, by induction on $m$, we have another interesting formula:
@@
It is straightforward to check that all elements in the six families listed in Proposition 1.7 are $P(2,1)$-invariants. By results in Section 3, it suffices to show that $\mathcal{B}_{m}(2,1)$ is a generating set for ${\mathcal{Q}}_{m}(3)^{P(2,1)}$.
Suppose $f$ is a nonzero polynomial in ${\mathcal{Q}}_{m}(3)^{P(2,1)}$, and we write $f$ as a polynomial in $x_{3}$:
$f=x_{3}^{a(q-1)}f_{a}+x_{3}^{(a-1)(q-1)}f_{a-1}+\ldots,$
We will prove by downward induction on the highest $x_{3}$-degree $a(q-1)$.
Observe that the coefficients $f_{i}$ belong to the space of invariants ${\mathcal{Q}}_{m}(2)^{G}$ in the variables $x_{1}$ and $x_{2}$. If $f_{a}$ is a scalar multiple of $x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$, then since $x_{3}^{a(q-1)}x_{1}^{q^{m}-1}x_{2}^{q^{m}-1}$ is already $P(2,1)$-invariant (which is a member of the 5th and 6th family in $\mathcal{B}_{m}(2,1)$), we can subtract it from $f$ to obtain a new one with strictly smaller $x_{3}$-degree.
Now let us assume that $f_{a}$, and hence all $f_{i}$, for degree reasons, are generated by $\delta_{2}(\Delta^{m}_{1})$ and $\Delta^{m}_{2}$. In particular, its $x_{1}$-degree (and hence $x_{2}$-degree as well, by symmetry) is strictly less than $q^{m}-1$. We claim that $a$ must be a multiple of $q$. Indeed, since $f$ is a $B$-invariant, from the list of generators given in 2.5, we see that $f$ must be a linear combination of polynomials of the form $x_{1}^{j_{1}(q-1)}\delta_{2}(Q_{2,1}^{j_{2}})$ or $x_{1}^{j_{1}(q-1)}Q_{2,1}^{j_{2}}Q_{3,2}^{j_{3}}$. In both cases, the highest $x_{3}$-degrees are divisible by $q$.
If $f_{a}$ contains $\delta_{2}(Q_{1,0}^{i_{1}})$ for some $i_{1}<[m]_{q}$, then we can subtract from $f$ an appropriate polynomial from the third or fourth family in the list 1.7. So we can assume that $f_{a}\in\Delta^{m}_{2}$. In this case, it follows that $x_{3}^{a(q-1}f_{a}$ is a nontrivial summand of a $B$-invariant polynomial of the form $x_{1}^{j_{1}(q-1)}Q_{2,1}^{j_{2}}Q_{3,2}^{j_{3}}$. In particular, $a$ must be divisible by $q^{2}$.
Finally, when $a(q-1)=iq^{2}(q-1)$ for some $i<[m-2]_{q}$ or $a(q-1)=q^{m}-q^{2}$, by subtracting from $f$ an appropriate linear combination of the first two families of $\mathcal{B}_{m}(2,1)$, we will obtain a $P(2,1)$-invariant polynomial of strictly smaller $x_{3}$-degree. The result follows by induction on the highest $x_{3}$-degree. ∎
@@
We can also prove that $\mathcal{B}_{m}(2,1)$ is linearly independent as follows. Consider these elements of $\mathcal{B}_{m}(2,1)$ as polynomials in $x_{3}$. The table below records the coefficient of the highest $x_{3}$-degree.
$\mathcal{B}_{m}(2,1)$ Highest $x_{3}$-degree Coefficient
@@
It suffices to show that $\mathcal{B}_{m}(1,2)$ is a generating set, and the proof is by downward induction on the lowest $x_{1}$-degree. Write